A piston-cylinder of gas has an initial pressure of 25 psi. The volume is increased to 2.5 times the initial volume. The temperature is increased from 290oC to 720oC. What is the final pressure. Be careful,.. SYW!... all of it
Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

P1V1/T1=P2V2/T2
25V1/290=P2*2.5V1/720
V1 cancel
25/290+2.5P2/720
720*25/290*2.5=P2
=24.83 psi
This makes intuitive sense since the temp has increased about 2.5 times and so has the volume, and they work against each other regarding volume,'

Answer by ikleyn(52772)

.
A piston-cylinder of gas has an initial pressure of 25 psi. The volume is increased to 2.5 times the initial volume.
The temperature is increased from 290oC to 720oC. What is the final pressure. Be careful,.. SYW!... all of it
~~~~~~~~~~~~~~~



Calculations by @Boreal are totally wrong due to different reasons,

one of which is that he does not convert Celsius temperatures to Kelvin temperatures.